OpenStudy (anonymous):

i need to find the slope of a line-parallel- to each given line. y=-x+5

OpenStudy (anonymous):

the lines is in the form y = mx + c where m = -1 in this equation which is nothing but the slope. And any line parallel to this line would have the same slope.

OpenStudy (anonymous):

Well, a parallel will have the same slope, so the slope will be -1. So, if you graph this line you know a point will be (0, 5) and (1, 4). By the pictures I attached, we can easily see two points which could be a parallel line. So, we do this using the point (1, 2) y-y1=-1(x-x1) y-2=-1(x-1) Now, put this in standard form, y=mx+b: y-2=-1x+2 +2 +2 y=-1x+4 The slope of this line will also be -1, just like the slope of line y=-x+5. Hope this helps. :o)

OpenStudy (anonymous):

oohh.. now i get it. Thank you.